# Differentiable Fuzzy Cosmic-Web for Field Level Inference

### v1

@@ 2.3 Gravity Solver
We model large-scale structure (LSS) formation with Augmented Lagrangian Perturbation Theory (ALPT) (Kitaura & Heß, 2013), which merges second-order LPT (2LPT) (Zel’dovich, 1970; Buchert, 1994; Bouchet et al., 1995; Catelan, 1995) for long-range tidal displacements and a spherical-collapse (SC) solution (Bernardeau, 1994; Neyrinck, 2016) for short-range dynamics. This is achieved by decomposing the particle displacement field as $\boldsymbol{\Psi}=\boldsymbol{\Psi}_{\rm L}+\boldsymbol{\Psi}_{\rm S},$ with $\boldsymbol{\Psi}_{\rm L}=\mathcal{K}(\boldsymbol{q},r_{s})\circ\boldsymbol{\Psi}_{\mathrm{2LPT}}$ and $\boldsymbol{\Psi}_{\rm S}=(1-\mathcal{K})\circ\boldsymbol{\Psi}_{\mathrm{SC}}$, with $\mathcal{K}$ a smoothing kernel of fixed radius $r_{s}=4$ $h^{-1}\,\mathrm{Mpc}$ following Kitaura & Heß (2013). In future work, we plan to sample this parameter within the Bayesian framework and consider nLPT for the long-range interaction.
This hybrid formulation retains analytic dependence on cosmological parameters while suppressing shell crossing. Coupled with empirical bias schemes such as BAM (Balaguera-Antolínez et al., 2018, 2020), ALPT reproduces $N$-body halo statistics down to $k\lesssim 0.4$$h\,\mathrm{Mpc}^{-1}$, the scale where most cosmological information resides (Hahn & Villaescusa-Navarro, 2021). An even higher accuracy is achieved with the HICOBIAN bias model considered in this work (see Sec. 2.7).
We note that ALPT is substantially more efficient in terms of both computational speed and memory usage compared to fast particle mesh (PM) methods (e.g., Tassev et al., 2013; Feng et al., 2016; Klypin & Prada, 2018), as shown in the comparison work by Blot et al. (2019), making it especially advantageous for scalable Bayesian inference frameworks. Moreover, ALPT can be extended via iteration (eALPT) to reach sub-Mpc accuracy (Kitaura et al., 2024), by indirectly emulating the effect of a viscous stress tensor in the equations of motion, thereby capturing aspects of nonlinear dynamics such as vorticity. Recent GPU-based PM methods have shown noteworthy advancements (e.g., Modi et al., 2021; Li et al., 2024). Alternative promising approaches to approximate structure formation are offered by emulators trained on large ensembles of simulations (Kodi Ramanah et al., 2020; Conceição et al., 2024). However, for the time being, we opt to rely on analytic solutions with explicit cosmological dependence.
@@ 2.4 Bias Model Framework
When employing an approximate gravity solver, the evolved density field should not be directly trusted at the particle level. Instead, it is reliable only on a coarse mesh with a resolution of a few Mpc, where the approximation remains accurate. As a result, the bias relation between matter and tracers is no longer deterministic–as in approaches that apply a halo finder directly to particle distributions–but must instead be modeled as an effective field-level bias (Schmittfull et al., 2019). In this framework, the luminous tracers of LSS exhibit a complex, nonlinear, nonlocal, and stochastic relationship with the underlying matter field (McDonald & Roy, 2009; Desjacques et al., 2018). Accurately capturing this relationship is essential for unbiased inference and requires flexible and physically motivated modelling approaches that operate at the field level.
We take advantage of a positive-definite, nonlocal bias modelling framework that employs local bias expansions within distinct morphological regions of the cosmic-web (Kitaura et al., 2022; Coloma-Nadal et al., 2024; Sinigaglia et al., 2024b).
2.5 Deterministic Local Bias Model
The bias relation between tracers and the underlying matter field has long been known to be nonlinear. One of the earliest approaches involved a local perturbative expansion of the tracer overdensity, introduced by Fry & Gaztanaga (1993):
@@ 2.5 Deterministic Local Bias Model
where $\delta_{\rm tr}$ denotes the tracer overdensity, $\delta$ the matter overdensity, and $b_{m}$ the bias coefficients. While conceptually straightforward, this expansion suffers from the drawback that it may yield unphysical negative tracer densities, when truncated at any order.
@@
which ensures positive-definiteness of the tracer field by construction. To linear order, this expansion corresponds to a power-law bias model, which has been particularly effective in modelling the Lyman-$\alpha$ forest. In this context, it leads to the fluctuating Gunn–Peterson approximation, where the optical depth $\tau$ is related to the matter density field by $\tau\propto(1+\delta)^{\alpha}$ (Bi & Davidsen, 1997). An analogous relationship was found for the functional dependency between the molecular gas temperature and dark matter density (Hui & Gnedin, 1997).
@@
Interestingly, a power-law bias model in Eulerian space can be related to a linear bias in Lagrangian space–i.e., in the coordinate system of the initial conditions–according to the continuity equation, which leads to a lognormal evolution of the density field (Coles & Jones, 1991). In this picture, a linear Lagrangian bias naturally evolves into a power-law bias at later times.
@@
In the context of galaxy bias, the polynomial perturbative expansion has historically been preferred. A foundational physical interpretation of bias was introduced even earlier by Kaiser (1984) through the excursion set model, where galaxies preferentially form in high-density peaks of the matter distribution. Building on this idea, Kitaura et al. (2014) proposed a hybrid model that combines the power-law bias of Cen & Ostriker (1993) with the threshold bias of Kaiser (1984) to model galaxy number counts as a function of the underlying dark matter field. This framework was later refined by Neyrinck et al. (2014), who replaced the hard threshold with an exponential cutoff. The resulting model captures the average bias behavior of halos with remarkable accuracy, particularly in low-density environments.
@@
Although the power-law and threshold bias models are degenerate with respect to two-point statistics–both amplifying power on large scales–this degeneracy can be broken using three-point statistics, which provide a means to discriminate between different biasing mechanisms (see Kitaura et al., 2015, also for the definitions of three-point statistics calculations used in this work). The hybrid model has proven especially useful in this regard and was employed to generate thousands of mock catalogs of luminous red galaxies for the final BOSS data release (see the PATCHY mocks, Kitaura et al., 2016b).
@@
Hitherto, the suppression of tracer densities has primarily focused on low-density regions, following the picture introduced by Kaiser. However, another important effect in the context of halo and galaxy formation is halo exclusion (Baldauf et al., 2013), which accounts for the saturation of tracer objects in highly overdense environments. This mechanism introduces a natural suppression of tracer abundance in high-density regions, complementing the low-density suppression from the original threshold bias picture (Coloma-Nadal et al., 2024). A similar concept has also been applied in modelling the Lyman-$\alpha$ forest, where saturation effects in flux absorption become significant in dense regions (Sinigaglia et al., 2024b).
@@
Motivated by these considerations, we model the expected tracer number density using a power-law component modulated by both a low-pass and a high-pass filter:
@@
$\bar{n}_{i}=C\left(1+\delta_{i}\right)^{\alpha}\exp\left[\left(\frac{1+\delta_{i}}{\rho_{\text{}}}\right)^{\epsilon_{\text{}}}\right]\exp\left[\left(\frac{1+\delta_{i}}{\rho^{\prime}_{\text{}}}\right)^{-\epsilon^{\prime}_{\text{}}}\right]\ ,$ (7)
where $C$ is a normalization constant defined such that $\sum_{i}\bar{n}_{i}=\sum_{i}n^{*}_{i}\equiv N_{\text{tr}}$, where $N_{\text{tr}}$ is the total number of tracers in the data. All parameters are positive-definite.
@@
We assume that the observed tracer count in each cell is a stochastic realization drawn from a discrete probability distribution $\mathcal{P}^{\text{tr}}_{i}$, characterized by a mean expected value $\bar{n}_{i}$ provided by the forward model.
@@ 2.6 Stochastic Bias Model
The coarse resolution inherent to the field-level framework necessitates the introduction of a stochastic component to the bias model (Dekel & Lahav, 1999; Sheth & Lemson, 1999), which is not explicitly present in deterministic high-resolution $N$-body simulations. Nonetheless, the lack of full phase-space information after shell-crossing also introduces an uncertainty which requires a probabilistic treatment in the likelihood.
@@
The most basic discretization assumes a Poisson likelihood, where the variance equals the mean tracer count. The complete integration of this likelihood beyond the second moment within the Bayesian inference framework has been studied for a long time (Kitaura & Enßlin, 2008; Kitaura et al., 2010). However, this assumption holds only in very limited regimes–typically for low tracer densities and in relatively homogeneous environments. As originally predicted by Peebles (1980), (anti-)correlations between tracers below the mesh resolution introduce deviations from pure Poisson statistics. These unresolved sub-grid correlations give rise to either sub- or super-Poissonian noise characteristics. To account for this, various probability distribution functions (PDFs) have been proposed in the literature (Saslaw & Hamilton, 1984; Sheth, 1995), most of which introduce additional degrees of freedom to model the overdispersion commonly observed in galaxy and halo distributions relative to the Poisson expectation (Somerville et al., 2001; Casas-Miranda et al., 2002). Pellejero-Ibañez et al. (2020) demonstrated that the Negative Binomial distribution can be parametrized to reproduce a wide range of functional dependencies in the over-dispersed deviations from Poisson statistics.
@@
We have implemented the Negative Binomial distribution to model $\mathcal{P}^{\text{tr}}_{i}$, thereby accounting for potential over-dispersion in the tracer counts. This distribution was first considered in the context of field-level bias in Kitaura et al. (2014) and later implemented in the Bayesian framework in Ata et al. (2015). Our implementation of the Negative Binomial distribution in this work relies on the Gamma–Poisson mixture modelling, in which the Poisson rate $\lambda_{i}$ is treated as a random variable drawn from a Gamma distribution. This hierarchical formulation is given by:
$\lambda_{i}$ $\sim\text{Gamma}(\beta,\bar{n}_{i}/\beta)\,,$ (8)
$n^{*}_{i}$ $\sim\text{Poisson}(\lambda_{i})\,,$ (9)
@@
where $\bar{n}_{i}$ denotes the expected tracer count in cell $i$, and $\beta$ is the over-dispersion parameter. In the limit $\beta\rightarrow\infty$, the Gamma distribution becomes sharply peaked, recovering the standard Poisson likelihood. In the Gamma distribution, the first argument is the shape parameter and the second argument is the scale parameter.. With this model, the (negative log) likelihood reads
$-\log\mathcal{L}(\boldsymbol{\nu},\boldsymbol{b},\beta)\equiv\sum_{i=1}^{N^{3}}-\log\mathcal{P}^{\text{tr}}_{i}(n^{*}_{i}|\mathcal{M}(\boldsymbol{\nu},\boldsymbol{b}),\beta)\ ,$ (10)
with $\boldsymbol{b}=(\alpha,\rho,\epsilon,\rho^{\prime},\epsilon^{\prime})$.
2.7 Nonlocal Bias and Cosmic Web Dependence
@@
In the previous sections, we introduced a local nonlinear bias model to describe the average bias relation and subsequently incorporated scatter to account for stochasticity. However, this stochastic component should be understood as a feature of the coarse-grained nature of the effective bias approach, without superseding deterministic contributions from nonlocal bias effects.
@@
Galaxy bias is shaped not only by the local matter density but also by the surrounding large-scale tidal environment and the small-scale curvature of the density field. These nonlocal dependencies stem from the intrinsic nature of gravitational collapse and galaxy formation. A general framework for incorporating such effects into bias expansions was introduced by McDonald & Roy (2009). In another particularly relevant study, Chan et al. (2012) connected long-range tidal tensor invariants to higher-order bias terms. Kitaura et al. (2022) later related the invariants of the Hessian of a field to the morphological classifications of the cosmic-web.
@@
These developments motivate extending bias models to explicitly include cosmic-web morphology (see Appendix B). Galaxies inhabiting different large-scale structures–voids, sheets, filaments, and knots–exhibit systematically different biasing behaviors, as shown in simulations and observations (e.g., Nuza et al., 2014; Filho et al., 2015).
@@
Several approaches exist for incorporating cosmic-web morphology. One method involves binning tidal field tensor invariants, which can result in a high-dimensional classification if multiple invariants and their dependence on local density are included (Kitaura et al., 2022). Alternatively, one can compress this information into four characteristic environments, sacrificing some modeling granularity for interpretability–this leads to the so-called $\Phi$-web classification, based on the large-scale tidal field (Hahn et al., 2007).
@@
To recover finer biasing distinctions, we introduce a hierarchical structure by embedding a second classification based on the Hessian of the density field–the $\delta$-web–within each $\Phi$-web region. This short-range nonlocal bias captures small-scale tidal effects, rooted in tidal torque theory (Heavens & Peacock, 1988). The result is the HICOBIAN model, which partitions the space into 16 regions (Coloma-Nadal et al., 2024). For each region, it applies a different local bias relation, determined by the specific bias parameters of model Eq. 7. This approach gains physical insight with computational efficiency and outperforms fine binning of the tidal field in terms of both precision and interpretability, as both short- and long-range nonlocal bias are incorporated.
@@
Given that gravitational evolution is inherently nonlocal, the shortcomings of ALPT relative to full gravity solvers can be effectively absorbed into nonlocal bias terms. This explains the excellent performance of more recent studies employing ALPT (e.g., Balaguera-Antolínez et al., 2023; Coloma-Nadal et al., 2024). This insight provides a foundation for extending the bias framework to capture deviations arising from alternative gravity models (García-Farieta et al., 2024), where characteristic bias parameters and higher-order statistics including Legendre multipole expansions would serve as indicators of such modified dynamics. However, it should be noted that even the most accurate gravity solvers, when down-sampled to a coarse mesh with a resolution of a few Mpc, introduce strong aliasing effects. These artifacts can also be effectively absorbed into nonlocal scale-dependent bias terms.
@@
In the context of this work, it is important to note that most cosmic-web classifiers rely on hard, non-differentiable thresholds (e.g., eigenvalue sign changes), which are incompatible with gradient-based inference techniques like Hamiltonian Monte Carlo. To address this, we adopt a fuzzy, differentiable classification scheme using sigmoid-based transitions. This enables smooth, physically motivated boundaries between web types and integrates seamlessly with our differentiable, field-level inference framework.
@@
Figure 2: Example of fuzzy cosmic-web classification with $N=256$ and $\Delta L=1.7\,h^{-1}\mathrm{Mpc}$. Top left: evolved dark matter density contrast field. Bottom left: “hard” $\Phi$-web classification. Right panels: fuzzy membership weights $p_{i}^{(\mathrm{V})}$, $p_{i}^{(\mathrm{S})}$, $p_{i}^{(\mathrm{F})}$, and $p_{i}^{(\mathrm{K})}$ for voids, sheets, filaments, and knots, respectively, as defined in Eq. 16.

### v3

@@ 2.3 Gravity solver
We modelled large-scale-structure (LSS) formation with Augmented Lagrangian perturbation theory (ALPT; Kitaura & Heß 2013), which merges second-order LPT (2LPT; Zel’dovich 1970; Buchert 1994; Bouchet et al. 1995; Catelan 1995) for long-range tidal displacements and a spherical-collapse (SC) solution (Bernardeau, 1994; Neyrinck, 2016) for short-range dynamics. This was achieved by decomposing the particle displacement field as implemented in the BRIDGE code:
$\boxed{\boldsymbol{\Psi}=\boldsymbol{\Psi}_{\rm L}+\boldsymbol{\Psi}_{\rm S}}\,$, (5)
with $\boldsymbol{\Psi}_{\rm L}=\mathcal{K}(\boldsymbol{q},r_{s})\circ\boldsymbol{\Psi}_{\mathrm{2LPT,}}$ $\boldsymbol{\Psi}_{\rm S}=(1-\mathcal{K})\circ\boldsymbol{\Psi}_{\mathrm{SC}}$, and $\mathcal{K}$ a smoothing kernel of fixed radius, $r_{s}=4$ $h^{-1}\,\mathrm{Mpc}$, following Kitaura & Heß (2013). In future work, we plan to sample this parameter within the Bayesian framework and consider nLPT for the long-range interaction.
This hybrid formulation retains analytic dependence on cosmological parameters while suppressing shell crossing. Coupled with empirical bias schemes such as the Bias Assignment Method (BAM). (Balaguera-Antolínez et al., 2018, 2020), ALPT reproduces $N$-body halo statistics down to $k\lesssim 0.4$$h\,\mathrm{Mpc}^{-1}$, the scale at which most cosmological information resides (Hahn & Villaescusa-Navarro, 2021). An even higher accuracy was achieved with the HICOBIAN bias model considered in this work (see Sect. 2.7).
We note that ALPT is substantially more efficient in terms of both computational speed and memory usage compared to fast particle mesh (PM) methods (e.g. Tassev et al., 2013; Feng et al., 2016; Klypin & Prada, 2018), as shown in the comparison work by Blot et al. (2019), making it especially advantageous for scalable Bayesian inference frameworks. Moreover, ALPT can be extended via iteration (eALPT) to reach sub-megaparsec accuracy (Kitaura et al., 2024) by indirectly emulating the effect of a viscous stress tensor in the equations of motion, thereby capturing aspects of non-linear dynamics such as vorticity. Recent GPU-based PM methods have shown noteworthy advancements (e.g. Modi et al., 2021; Li et al., 2024). Alternative promising approaches to approximate structure formation are offered by emulators trained on large ensembles of simulations (Kodi Ramanah et al., 2020; Conceição et al., 2024). However, for the time being, we opt to rely on analytic solutions with explicit cosmological dependence.
@@ 2.4 Bias model framework
When employing an approximate gravity solver, the evolved density field should not be directly trusted at the particle level. Instead, it is only reliable on a coarse mesh with a resolution of a few megaparsecs, where the approximation remains accurate. As a result, the bias relation between matter and tracers is no longer deterministic – as in approaches that apply a halo finder directly to particle distributions – it must be modelled as an effective field-level bias (Schmittfull et al., 2019). In this framework, the luminous tracers of LSS exhibit a complex, non-linear, non-local, and stochastic relationship with the underlying matter field (McDonald & Roy, 2009; Desjacques et al., 2018). Accurately capturing this relationship is essential for unbiased inference and requires flexible and physically motivated modelling approaches that operate at the field level.
We took advantage of a positive, non-local-bias modelling framework that employs local-bias expansions within distinct morphological regions of the cosmic web (Kitaura et al., 2022; Coloma-Nadal et al., 2024; Sinigaglia et al., 2024b).
2.5 Deterministic local-bias model
The bias relation between tracers and the underlying matter field has long been known to be non-linear. One of the earliest approaches involved a local perturbative expansion of the tracer overdensity, introduced by Fry & Gaztanaga (1993):
@@ 2.5 Deterministic local-bias model
where $\delta_{\rm tr}$ denotes the tracer overdensity, $\delta$ the matter overdensity, and $b_{m}$ the bias coefficients. While conceptually straightforward, this expansion suffers from the drawback that it may yield unphysical negative tracer densities when truncated at any order.
@@
which ensures positivity of the tracer field by construction. To a linear order, this expansion corresponds to a power-law bias model, which has been particularly effective in modelling the Lyman-$\alpha$ forest. In this context, it leads to the fluctuating Gunn–Peterson approximation, where the optical depth, $\tau,$ is related to the matter density field by $\tau\propto(1+\delta)^{\alpha}$ (Bi & Davidsen, 1997). An analogous relationship was found for the functional dependency between the molecular gas temperature and dark-matter density (Hui & Gnedin, 1997).
@@
Interestingly, a power-law bias model in Eulerian space can be related to a linear bias in Lagrangian space – i.e. in the coordinate system of the initial conditions – according to the continuity equation, which leads to a log-normal evolution of the density field (Coles & Jones, 1991). In this picture, a linear Lagrangian bias naturally evolves into a power-law bias at later times.
@@
In the context of galaxy bias, the polynomial perturbative expansion has historically been preferred. A foundational physical interpretation of bias was introduced even earlier by Kaiser (1984) through the excursion set model, where galaxies preferentially form in high-density peaks of the matter distribution. Building on this idea, Kitaura et al. (2014) proposed a hybrid model that combines the power-law bias of Cen & Ostriker (1993) with the threshold bias of Kaiser (1984) to model galaxy number counts as a function of the underlying dark-matter field. This framework was later refined by Neyrinck et al. (2014), the authors of which replaced the hard threshold with an exponential cutoff. The resulting model captures the average bias behaviour of halos with remarkable accuracy, particularly in low-density environments.
@@
Although the power-law and threshold bias models are degenerate with respect to two-point statistics – both amplifying power on large scales – this degeneracy can be broken using three-point statistics, which provide a means to distinguish between different biasing mechanisms (see also Kitaura et al. 2015 for the definitions of three-point statistics calculations used in this work). The hybrid model has proven especially useful in this regard and was employed to generate thousands of mock catalogues of luminous red galaxies for the final BOSS data release (see the PATCHY mocks in Kitaura et al., 2016b).
@@
To date, the suppression of tracer densities has primarily focused on low-density regions, following the picture introduced by Kaiser. However, another important effect in the context of halo and galaxy formation is halo exclusion (Baldauf et al., 2013), which accounts for the saturation of tracer objects in highly overdense environments. This mechanism introduces a natural suppression of tracer abundance in high-density regions, complementing the low-density suppression from the original threshold bias picture (Coloma-Nadal et al., 2024). A similar concept has also been applied in modelling the Lyman-$\alpha$ forest, where saturation effects in flux absorption become significant in dense regions (Sinigaglia et al., 2024b).
@@
Motivated by these considerations, we modelled and implemented the expected tracer number density in the BRIDGE code using a power-law component modulated by both a low-pass and a high-pass filter:
@@
$\boxed{\bar{n}_{i}=C\left(1+\delta_{i}\right)^{\alpha}\exp\left[-\left(\frac{1+\delta_{i}}{\rho_{\text{}}}\right)^{\epsilon_{\text{}}}\right]\exp\left[-\left(\frac{1+\delta_{i}}{\rho^{\prime}_{\text{}}}\right)^{-\epsilon^{\prime}_{\text{}}}\right]}\ ,$ (8)
@@
where $C$ is a normalisation constant defined such that $\sum_{i}\bar{n}_{i}=\sum_{i}n^{*}_{i}\equiv N_{\text{tr}}$, where $N_{\text{tr}}$ is the total number of tracers in the data. All parameters are positive. We assume that the observed tracer count in each cell is a stochastic realisation drawn from a discrete probability distribution, $\mathcal{P}^{\text{tr}}_{i}$, characterised by a mean expected value, $\bar{n}_{i}$, provided by the forward model.
@@ 2.6 Stochastic bias model
The coarse resolution inherent to the field-level framework necessitates the introduction of a stochastic component to the bias model (Dekel & Lahav, 1999; Sheth & Lemson, 1999), which is not explicitly present in deterministic high-resolution $N$-body simulations. Nonetheless, the lack of full phase-space information after shell-crossing also introduces an uncertainty that requires a probabilistic treatment in the likelihood.
@@
The most basic discretisation assumes a Poisson likelihood, where the variance equals the mean tracer count. The complete integration of this likelihood beyond the second moment within the Bayesian inference framework has been studied for a long time (Kitaura & Enßlin, 2008; Kitaura et al., 2010). However, this assumption holds only in very limited regimes – typically for low tracer densities and in relatively homogeneous environments. As originally predicted by Peebles (1980), (anti-)correlations between tracers below the mesh resolution introduce deviations from pure Poisson statistics. These unresolved sub-grid correlations give rise to either sub- or super-Poissonian noise characteristics. To account for this, various probability distribution functions (PDFs) have been proposed in the literature (Saslaw & Hamilton, 1984; Sheth, 1995), most of which introduce additional degrees of freedom to model the over-dispersion commonly observed in galaxy and halo distributions relative to the Poisson expectation (Somerville et al., 2001; Casas-Miranda et al., 2002). Pellejero-Ibañez et al. (2020) demonstrated that the negative binomial distribution can be parametrised to reproduce a wide range of functional dependencies in the over-dispersed deviations from Poisson statistics.
@@
We implemented the negative binomial distribution to model $\mathcal{P}^{\text{tr}}_{i}$, thereby accounting for potential over-dispersion in the tracer counts. This distribution was first considered in the context of field-level bias in Kitaura et al. (2014) and later implemented in the Bayesian framework in Ata et al. (2015). Our implementation of the negative binomial distribution in this work relies on the gamma–Poisson mixture modelling, in which the Poisson rate, $\lambda_{i},$ is treated as a random variable drawn from a gamma distribution. The corresponding hierarchical formulation implemented in the BRIDGE code is given by
$\boxed{\lambda_{i}\sim\text{Gamma}(\beta,\bar{n}_{i}/\beta)}$ (9)
$\boxed{n^{*}_{i}\sim\text{Poisson}(\lambda_{i})}$ $\,$, (10)
@@
where $\bar{n}_{i}$ denotes the expected tracer count in cell $i$, and $\beta$ is the over-dispersion parameter. In the limit $\beta\rightarrow\infty$, the gamma distribution becomes sharply peaked, recovering the standard Poisson likelihood. In the gamma distribution, the first argument is the shape parameter and the second argument is the scale parameter. With this model, the (negative log) likelihood reads
$\log\mathcal{L}(\boldsymbol{\nu},\boldsymbol{b},\beta)\equiv\sum_{i=1}^{N^{3}}-\log\mathcal{P}^{\text{tr}}_{i}(n^{*}_{i}|\mathcal{M}(\boldsymbol{\nu},\boldsymbol{b}),\beta)\ ,$ (11)
with $\boldsymbol{b}=(\alpha,\rho,\epsilon,\rho^{\prime},\epsilon^{\prime})$ from Eq. 8. While sub-Poissonian behaviour is indirectly captured by the high-density damping exponential factor, an explicit treatment can be implemented using the Conway–Maxwell–Poisson distribution (Daly & Gaunt, 2016) or extensions of the binomial distribution (Vos-Ginés et al., 2024).
2.7 Non-local bias and cosmic-web dependence
le
@@ 2.7 Non-local bias and cosmic-web dependence
In the previous sections, we introduced a local non-linear-bias model to describe the average bias relation and subsequently incorporated scatter to account for stochasticity. However, this stochastic component should be understood as a feature of the coarse-grained nature of the effective bias approach, without superseding deterministic contributions from non-local-bias effects.
@@
Galaxy bias is shaped not only by the local matter density, but also by the surrounding large-scale tidal environment and the small-scale curvature of the density field. These non-local dependencies stem from the intrinsic nature of gravitational collapse and galaxy formation. A general framework for incorporating such effects into bias expansions was introduced by McDonald & Roy (2009). In another particularly relevant study, Chan et al. (2012) connected long-range tidal tensor invariants to higher order bias terms. Kitaura et al. (2022) later related the invariants of the Hessian of a field to the morphological classifications of the cosmic web.
@@
These developments motivated the extension of bias models to explicitly include cosmic-web morphology (see Appendix B). Galaxies inhabiting different large-scale structures, including voids, sheets, filaments, and knots, exhibit systematically different biasing behaviours, as shown in simulations and observations (e.g. Nuza et al., 2014; Filho et al., 2015).
@@
Several approaches exist for incorporating cosmic-web morphology. One method involves binning tidal field tensor invariants, which can result in a high-dimensional classification if multiple invariants and their dependence on local density are included (Kitaura et al., 2022). Alternatively, one can compress this information into four characteristic environments, sacrificing some modelling granularity for interpretability – this leads to the so-called $\Phi$-web classification, based on the large-scale tidal field (Hahn et al., 2007).
@@
To recover finer biasing distinctions, we introduced a hierarchical structure by embedding a second classification based on the Hessian of the density field (the $\delta$-web) within each $\Phi$-web region. This short-range non-local bias captures small-scale tidal effects rooted in tidal torque theory (Heavens & Peacock, 1988). The result is the HICOBIAN model, which partitions the space into 16 regions (Coloma-Nadal et al., 2024). For each region, it applies a different local-bias relation, which is determined by the specific bias parameters of model Eq. 8. This approach gains physical insight with computational efficiency and outperforms fine binning of the tidal field in terms of both precision and interpretability, as both short- and long-range non-local biases are incorporated.
@@
Given that gravitational evolution is inherently non-local, the shortcomings of ALPT relative to full gravity solvers can be effectively absorbed into non-local-bias terms. This explains the excellent performance of more recent studies employing ALPT (e.g. Balaguera-Antolínez et al., 2023; Coloma-Nadal et al., 2024). This insight provides a foundation for extending the bias framework to capture deviations arising from alternative gravity models (García-Farieta et al., 2024), where characteristic bias parameters and higher order statistics including Legendre multipole expansions would serve as indicators of such modified dynamics.
@@
In the context of this work, it is important to note that most cosmic-web classifiers rely on hard, non-differentiable thresholds (e.g. eigenvalue sign changes), which are incompatible with gradient-based inference techniques such as HMC sampling. To address this, we adopted a fuzzy, differentiable classification scheme using sigmoid-based transitions. This enables smooth, physically motivated boundaries between web types and integrates seamlessly with our differentiable, field-level inference framework.
@@
Figure 2: Example of fuzzy cosmic-web classification with $N=256$ and $\Delta L=1.7\,h^{-1}\mathrm{Mpc}$. Top left: Evolved dark-matter density contrast field. Bottom left: ‘Hard’ $\Phi$-web classification. Right panels: Fuzzy membership weights, $p_{i}^{(\mathrm{V})}$, $p_{i}^{(\mathrm{S})}$, $p_{i}^{(\mathrm{F})}$, and $p_{i}^{(\mathrm{K})}$, for voids, sheets, filaments, and knots, respectively, as defined in Eq. 17.